# A Neurosymbolic Fast and Slow Architecture for Graph Coloring

### v1

@@ Abstract
Constraint Satisfaction Problems (CSPs) present significant challenges to artificial intelligence due to their intricate constraints and the necessity for precise solutions. Existing symbolic solvers are often slow, and prior research has shown that Large Language Models (LLMs) alone struggle with CSPs because of their complexity. To bridge this gap, we build upon the existing SOFAI architecture (or SOFAI-v1), which adapts Daniel Kahneman’s “Thinking, Fast and Slow” cognitive model to AI. Our enhanced architecture, SOFAI-v2, integrates refined metacognitive governance mechanisms to improve adaptability across complex domains, specifically tailored here for solving CSPs like graph coloring. SOFAI-v2 combines a fast System 1 (S1) based on LLMs with a deliberative System 2 (S2) governed by a metacognition module. S1’s initial solutions, often limited by non-adherence to constraints, are enhanced through metacognitive governance, which provides targeted feedback and examples to adapt S1 to CSP requirements. If S1 fails to solve the problem, metacognition strategically invokes S2, ensuring accurate and reliable solutions. With empirical results, we show that SOFAI-v2 for graph coloring problems achieves an 16.98% increased success rate and is 32.42% faster than symbolic solvers.
Introduction
@@
Constraint Satisfaction Problems are a core challenge in artificial intelligence (AI) due to their demand for correctness adhering to strict constraints in static (Kumar 1992) or uncertain, dynamic environments (Verfaillie and Jussien 2005). These problems are prominent in applications such as scheduling, boolean satisfiability, resource allocation, temporal reasoning, and planning (Welsh and Powell 1967; Chaitin 1982). Traditional approaches to solve CSPs typically rely on constraint propagation or search (Haralick and Elliott 1980). The symbolic approaches are known for their accuracy but often struggle with scalability as the complexity of CSPs increases, hindered by high computational overhead and slow processing times (Dechter 2003; Rossi, Van Beek, and Walsh 2006).
@@
Recent advancements in LLMs have shown promise in rapidly processing complex information (Ruoss et al. 2024), but they are fundamentally limited in solving sequential decision-making tasks like planning (Valmeekam et al. 2022), and reasoning (Stechly, Valmeekam, and Kambhampati 2024). Due to their probabilistic, retrieval-based nature, LLMs approximate solutions rather than generating definitive answers, often leading to partial or inconsistent adherence to constraints (Jiang et al. 2023). Even with state-of-the-art prompting strategies and self-verification techniques, LLMs fall short in handling the strict requirements of CSPs, making their outputs unreliable for applications demanding high precision.
@@
The disparity between the precision of traditional symbolic methods and the adaptability and speed of LLMs underscores the need for a new paradigm in tackling complex tasks like graph coloring (Jensen and Toft 2011), a canonical CSP. Drawing inspiration from Daniel Kahneman’s cognitive theory of “Thinking, Fast and Slow” (Kahneman 2011) and its adaptation to AI as SOFAI-v1 (Booch et al. 2021), we introduce SOFAI-v2 which extends the earlier architecture by refining metacognitive mechanisms that oversee a fast, experience-based S1, powered by an LLM, and a slow, deliberative S2 using the DSATUR algorithm for graph coloring. This governance mechanism continuously monitors and enhances S1’s outputs, offering targeted feedback and constraint adherence examples; when S1’s solutions fall short, S2 is strategically invoked to guarantee accuracy. In this setup, S1 swiftly produces initial solutions, while S2 provides a reliable fallback for achieving precise constraint satisfaction. Our contributions are as follows:
@@
  • We introduce SOFAI-v2, a neurosymbolic fast and slow architecture designed to enhance the performance of CSP solvers on graph coloring problems through adaptive metacognitive governance.
@@
  • We develop a comprehensive benchmark suite of graph coloring problems featuring variations in graph size, edge probability, and a mix of solvable and unsolvable instances to test solver adaptability.
@@
  • We conduct extensive empirical evaluations of various solver configurations for graph coloring, demonstrating the advantages of SOFAI-v2 in terms of success rate, time efficiency, and adaptability.
@@
Background
Constraint Satisfaction Problem
CSP is defined as a tuple $(X,D,C)$, where:
  • $X=\{x_{1},x_{2},\dots,x_{n}\}$ is a finite set of variables.
  • $D=\{D_{1},D_{2},\dots,D_{n}\}$ is a set of domains, with each $D_{i}$ representing the possible values for variable $x_{i}$.
  • $C=\{c_{1},c_{2},\dots,c_{m}\}$ is a set of constraints, with each constraint $c_{j}$ defining a subset of allowed assignments for a set of variables involved in the constraint.
A solution to a CSP is an assignment of values to all variables in $X$ such that every constraint in $C$ is satisfied. CSPs are used extensively in scheduling, resource allocation, and combinatorial optimization, where satisfying multiple constraints simultaneously is crucial.
Graph Coloring Problem
@@
The graph coloring problem is a specific type of CSP, where the goal is to assign a color to each node of a graph such that no two adjacent nodes share the same color. Formally, it can be defined as a tuple $(G,C,k)$, where:
  • $G=(V,E)$ is an undirected graph, with $V=\{v_{1},v_{2},\dots,v_{n}\}$ representing the set of vertices (or nodes) and $E=\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$ representing the set of edges.
  • $C=\{c_{1},c_{2},\dots,c_{k}\}$ is a set of $k$ colors.
  • The objective is to find a function $f:V\rightarrow C$ such that for any edge $(v_{i},v_{j})\in E$, $f(v_{i})\neq f(v_{j})$.
We use the DIMACS representation (Johnson and Trick 1996) to describe the graph coloring problem, and its corresponding visualization is shown in Figure 1.
1 c Example of a graph coloring problem in DIMACS format
2 p edge 5 5
3 c edges
4 e A B
5 e A C
6 e B C
7 e C D
8 e D E
(a) DIMACS Format of Graph Coloring Problem
(b) Visualization of the Graph
Figure 1: Graph Coloring Problem in DIMACS Format and Its Visualization
Degree of Saturation Algorithm with Backtracking
The Degree of Saturation, or DSATUR algorithm is a heuristic for approaching optimal coloring by assigning colors to vertices based on their saturation degree—the number of distinct colors already assigned to adjacent vertices (Brélaz 1979). While the algorithm by itself is not guaranteed to find an optimal solution, the addition of backtracking allows it to systematically explore alternative color assignments, ensuring optimal coloring is achieved.
Given the graph coloring problem $(G,C,k)$, where $G=(V,E)$ is an undirected graph, $C=\{c_{1},c_{2},\dots,c_{k}\}$ is a set of colors, and $f:V\rightarrow C$ is the coloring function, the algorithm proceeds as follows:
  1. 1.
    Initialization: Set $f(v)=\emptyset$ for all $v\in V$, marking all vertices as uncolored.
  2. 2.
    Saturation Degree Calculation: For each uncolored vertex $v\in V$, calculate its saturation degree, $d_{sat}(v)$, defined as the number of distinct colors assigned to its adjacent vertices: $\{f(u)\mid(v,u)\in E,f(u)\neq\emptyset\}$.
  3. 3.
    Vertex Selection: Choose the uncolored vertex $v$ with the highest saturation degree $d_{sat}(v)$. In the case of a tie, select the vertex with the highest degree $d(v)$, which is the number of edges incident to $v$.
  4. 4.
    Color Assignment: Assign the smallest feasible color $c\in C$ to $v$, ensuring that $f(v)\neq f(u)$ for all adjacent vertices $(v,u)\in E$.
  5. 5.
    Backtracking (to Ensure Optimality): If no valid color can be assigned without violating constraints, initiate backtracking. Backtracking systematically undoes recent assignments and explores alternate color choices. This recursive search allows the algorithm to revisit decisions, thereby exploring all possible color assignments and guaranteeing that an optimal solution—one using the minimal number of colors—is achieved.
  6. 6.
    Repeat: Continue steps 2-5 until all vertices are colored or no valid coloring exists.
The algorithm finds a minimal coloring assignment by prioritizing vertices with the highest constraints and employing backtracking to explore alternatives. Without backtracking, the heuristic alone may fall short of optimality; with backtracking, however, the algorithm can reach an optimal solution.
Related Works
@@
Figure 2: SOFAI-v2 architecture with real-time metacognitive governance
@@
This section reviews the evolution of solvers for graph coloring, the application of LLMs in sequential decision-making tasks, and the emergence of neurosymbolic approaches for enhancing decision-making processes.
Solvers for CSP
@@
Traditional symbolic solvers have been extensively developed for this problem, leveraging algorithms that systematically explore the search space. Early works like Brélaz (1979) introduced efficient heuristics such as the DSATUR algorithm, significantly improving practical performance. Chaitin (1982) applied graph coloring to compiler optimization, demonstrating the versatility of symbolic methods. Golumbic (2004) provided a comprehensive treatment of graph algorithms, including coloring techniques. Despite their accuracy, symbolic approaches often struggle with scalability as CSP complexity increases, hindered by high computational overhead and slow processing times (Dechter 2003). Techniques like constraint propagation (Mackworth 1977) and backtracking algorithms (Bitner and Reingold 1975) have been employed to enhance efficiency but are limited by the exponential growth of the search space. To address these limitations, neural solvers have been explored, utilizing machine learning techniques to approximate solutions for graph coloring. For instance, Li, Chen, and Koltun (2018) proposed using graph neural networks to tackle combinatorial optimization problems. Similarly, Khalil et al. (2017) introduced a framework that learns heuristics directly from data, improving scalability. However, neural approaches often face challenges in strictly adhering to the constraints inherent in CSPs.
LLMs for Reasoning Tasks
@@
Recent developments in LLMs have opened new avenues for addressing sequential decision-making tasks, including CSPs. Various prompt engineering techniques have been proposed to enhance the reasoning capabilities of LLMs (Brown 2020; Wei et al. 2022; Kojima et al. 2022; Nye et al. 2021; Wang et al. 2022). Han et al. (2023) demonstrated that LLMs can arrive at correct answers for graph-based problems through iterative prompting. However, the reasoning capabilities of LLMs have been found to be fundamentally approximate retrieval in nature (Kambhampati 2024), limiting their effectiveness in precise reasoning tasks. This limitation is evidenced by studies showing LLMs’ inability to solve planning problems (Valmeekam et al. 2022) and graph coloring problems (Stechly, Valmeekam, and Kambhampati 2024), where strict adherence to constraints is required (Jiang et al. 2023).
Neurosymbolic Approaches for Reasoning Tasks
@@
Neurosymbolic approaches have gained considerable attention in addressing complex sequential decision-making tasks such as automated planning (Fabiano et al. 2023), constrained grid navigation (Ganapini et al. 2022), and puzzle solving (Lin et al. 2024). These studies demonstrate that integrating neural methods, such as LLMs, with symbolic solvers enhances performance on sequential decision-making tasks. To the best of our knowledge, neurosymbolic techniques have not been explored in the context of graph coloring problems.
The SOFAI Architecture
@@
The SOFAI architecture, also referred to as SOFAI-v1 here (Booch et al. 2021), leverages a rule-based metacognitive control mechanism that dynamically chooses between S1 and S2 solvers based on a confidence threshold associated with S1’s outputs. S1 solvers are experience-based (and usually data-driven) solvers, while S2 solvers are deliberative (and usually symbolic and rule-based) solvers. SOFAI-v1 has been instantiated to planning (Fabiano et al. 2023) and constrained-grid navigation (Ganapini et al. 2022), showing improvement on the performance compared to either symbolic or data-driven solvers for the same class of problems.
@@
In SOFAI-v1, the metacognitive component only chooses between S1 and S2 solvers, but does not exploit the possible collaboration between these two kinds of solvers, nor does it provide feedback to the solvers if they cannot solve the given problem instance. In this paper we consider a generalized version of this architecture, called SOFAI-v2, shown in Figure 2, where the metacognitive component (called MC) can provide feedback and examples to a failing S1 solver (that for this paper is an LLM), calling the S1 solver more than once, until either a correct solution is returned or a maximum number of iterations is reached, which triggers the activation of an S2 solver.
@@
Formally, let $f_{\text{S1}}(x,M)$ denote the solution produced by S1 for a problem instance $x$ with access to an episodic memory $M$. Here, $M$ consists of previously encountered problem-solution pairs generated by SOFAI-v2, i.e., pairs $(x_{j},f_{\text{SOFAI-v2}}(x_{j}))$, where $f_{\text{SOFAI-v2}}(x_{j})$ represents the final solution that could have been generated by either S1 or S2.
@@
where $x_{j}\in M$ and $\alpha$ is a predefined similarity threshold. This subset $M_{x}$ provides S1 with additional contextual information from past solutions, enabling it to leverage the SOFAI-v2’s previous outputs for experience-based adaptation.
Let $C(f_{\text{S1}}(x,M_{x}))$ denote the confidence score assigned by MC to the solution $f_{\text{S1}}(x,M_{x})$, produced by S1 using both the current problem instance $x$ and the retrieved memory subset $M_{x}$. If $C(f_{\text{S1}}(x,M_{x}))\geq\theta_{\text{S1}}$, where $\theta_{\text{S1}}$ is a predefined confidence threshold, the solution is accepted as final. Otherwise, MC provides feedback $\mathcal{F}(f_{\text{S1}}(x,M_{x}))$ or generates a simplified example $\mathcal{E}(x)$, where $\mathcal{E}(x)$ is a reduced version of $x$, and calls the S1 solver again with the additional information included in the feedback and the examples.
MC also monitors the trend in S1’s confidence scores over $k$ iterations to evaluate the improvement. Let $C_{i}=C(f_{\text{S1}}(x_{i},M_{x_{i}}))$ represent the confidence score after the $i$-th iteration on similar or modified instances. An improvement trend is identified if:
$\forall i\in\{1,\dots,k-1\},\quad C_{i+1}>C_{i}.$
If this trend condition is not satisfied after $k$ iterations, MC gives the problem instance to S2, invoking $f_{\text{S2}}(x)$ as a fallback to solve it. This ensures robustness, particularly in cases where S1’s adaptation is insufficient to meet confidence requirements, even with feedback, examples, and episodic memory support.
Experimental Setup and Results
This section details the methodology and results of experiments designed to evaluate the performance of various solver configurations on graph coloring problems. The experiments focus on efficiency and success rate, incorporating a systematic approach to data generation, problem classification, and solver configuration.
Data Generation and Problem Classification

### v2

@@ Abstract
Constraint Satisfaction Problems (CSPs) present significant challenges to artificial intelligence due to their intricate constraints and the necessity for precise solutions. Existing symbolic solvers are often slow, and prior research has shown that Large Language Models (LLMs) alone struggle with CSPs because of their complexity. To bridge this gap, we build upon the existing SOFAI architecture (SOFAI-v1), which adapts Daniel Kahneman’s “Thinking, Fast and Slow” cognitive model to AI. Our enhanced architecture, SOFAI-v2, integrates refined metacognitive governance mechanisms to improve adaptability across complex domains, specifically tailored here for solving the graph coloring problem, a specific type of CSP. SOFAI-v2 combines a fast System 1 (S1), leveraging LLMs, with a deliberative System 2 (S2), governed by a metacognition module. S1’s initial solutions, often limited by constraint adherence issues, are improved through targeted feedback and examples from metacognition, aligning S1 more closely with CSP requirements. If S1 fails to resolve the problem, metacognition strategically invokes S2, ensuring accurate and reliable solutions. Our empirical results demonstrate that SOFAI-v2 achieves a 10.5% higher success rate and is up to 30% faster than a traditional symbolic solver in solving graph coloring problems.
vedant@email.sc.edu vishalp@mailbox.sc.edu biplav.s@sc.edu francesca.rossi2@ibm.com
1University of South Carolina, Columbia, SC, USA.
2IBM Research, NY, USA.
1 Introduction
@@
Constraint Satisfaction Problems are a core challenge in artificial intelligence (AI) due to their demand for correctness adhering to strict constraints in static (Kumar, 1992) or uncertain, dynamic environments (Verfaillie & Jussien, 2005). These problems are prominent in applications such as scheduling, boolean satisfiability, resource allocation, temporal reasoning, and planning (Welsh & Powell, 1967; Chaitin, 1982). Traditional approaches to solve CSPs typically rely on constraint propagation or search (Haralick & Elliott, 1980). The symbolic approaches are known for their accuracy but often struggle with scalability as the complexity of CSPs increases, hindered by high computational overhead and slow processing times (Dechter, 2003; Rossi et al., 2006).
@@
Recent advancements in LLMs have shown promise in rapidly processing complex information (Ruoss et al., 2024), but they are fundamentally limited in solving sequential decision-making tasks like planning (Valmeekam et al., 2022), and reasoning (Stechly et al., 2024). Due to their probabilistic, retrieval-based nature, LLMs approximate solutions rather than generating definitive answers, often leading to partial or inconsistent adherence to constraints (Jiang et al., 2023). Even with state-of-the-art prompting strategies and self-verification techniques, LLMs fall short in handling strict requirements of CSPs, making their outputs unreliable for applications demanding high precision. We study feasibility under fixed resource limits in the graph coloring decision problem, a core constraint satisfaction setting that supports clear verification and iterative guidance.
@@
The disparity between the precision of traditional symbolic methods and the adaptability and speed of LLMs underscores the need for a new paradigm in tackling complex tasks like graph coloring (Jensen & Toft, 2011), a canonical CSP. Drawing inspiration from Daniel Kahneman’s cognitive theory of “Thinking, Fast and Slow” (Kahneman, 2011) and its adaptation to AI as SOFAI-v1 (Fabiano et al., 2025; Booch et al., 2021), we introduce SOFAI-v2 which extends the earlier architecture by refining metacognitive mechanisms that oversee a fast, experience-based S1, powered by an LLM, and a slow, deliberative S2 using the DSATUR algorithm for graph coloring. This governance mechanism continuously monitors and enhances S1’s outputs, offering targeted feedback and constraint adherence examples; when S1’s solutions fall short, S2 is strategically invoked to guarantee accuracy. In this setup, S1 swiftly produces initial solutions, while S2 provides a reliable fallback for achieving precise constraint satisfaction. This work studies problem solving with a fast S1, a deliberate S2, and an MC that monitors and guides iterations, with memory for reuse. Beyond the S1/S2 framing, SOFAI-v2 is positioned as a neurosymbolic cognitive architecture: S1 (neural) proposes, symbolic constraints verify, and MC enacts control. This aligns with computational accounts of automatic versus controlled processing and supervisory control (e.g., contention scheduling and supervisory attention) (Norman & Shallice, 1986; Cooper & Shallice, 2000), providing a bridge from cognitive theory to an AI system.
The implemented system targets the graph coloring decision task and evaluates success rate, time, and iteration behavior. Our contributions are as follows:
@@
  • Introduced SOFAI-v2, a neurosymbolic fast and slow architecture to enhance the performance of CSP solvers on graph coloring problems through adaptive metacognitive governance.
@@
  • Created a comprehensive benchmark of graph coloring problems, featuring variations in graph size, edge probability, and solvable and unsolvable instances to test adaptability.
@@
  • Conducted extensive empirical evaluations of various solver configurations for graph coloring, demonstrating advantages of SOFAI-v2 in terms of success rate, time efficiency.
@@
2 Background
2.1 Graph Coloring Decision Problem
@@
The graph coloring decision problem is a specific type of constraint satisfaction problem (CSP), where the objective is to determine whether a given graph can be colored using at most $k$ colors such that no two adjacent nodes share the same color. Formally, the problem is defined as follows:
  • An undirected, unweighted graph $G=(V,E)$
  • A positive integer $k$ representing the maximum number of colors
  • Determine whether there exists a function $f:V\rightarrow\{1,2,\dots,k\}$ such that for every edge $(u,v)\in E$, $f(u)\neq f(v)$.
Unlike the classical graph coloring problem, which aims to minimize the number of colors used (an optimization problem), the decision version focuses solely on satisfiability: whether a valid coloring exists for a given $k$. We use the DIMACS representation (Johnson & Trick, 1996) to describe the graph coloring problem. The decision problem asks a clear yes/no question for a fixed $k$. This matches many practical settings where a fixed resource limit is given and the goal is feasibility, e.g., schedules with a fixed number of time slots, register allocation with a fixed number of registers, or frequency assignment with a fixed set of bands. It provides MC with a simple accept/reject signal and makes it straightforward to generate targeted feedback when constraints fail, which we interpret as a cue for supervisory control in the architecture (Norman & Shallice, 1986; Cooper & Shallice, 2000).
2.2 Degree of Saturation Algorithm with Backtracking
The Degree of Saturation (DSATUR) algorithm assigns colors to graph vertices based on their saturation degree, prioritizing vertices that are the most constrained—those having adjacent vertices with the greatest variety of colors already assigned (Brélaz, 1979). DSATUR remains competitive for graph coloring due to its heuristic efficiency, especially for sparse and mid-sized graphs, and its direct operation on graph structures without requiring transformations typical of SAT solvers. Recent studies (Yekezare et al., 2024; Schidler & Szeider, 2023) highlight its effectiveness, making it a suitable choice for our symbolic solver. However, the traditional DSATUR algorithm targets the optimization version of the graph coloring problem, minimizing the number of colors. For our work, we adapt DSATUR into a satisfiability solver by constraining it with a fixed maximum number of colors. A detailed description of the modified algorithm is provided in Supplementary Material Section 7. Unlike other solvers, DSATUR can be straightforwardly modified for the decision version of the problem, making it a practical and adaptable choice for our symbolic backend. In SOFAI-v2, DSATUR with backtracking plays the role of slow, controlled processing that is selectively invoked by MC under conflict or low confidence, consistent with supervisory control accounts in cognitive systems (Norman & Shallice, 1986; Cooper & Shallice, 2000).
3 Related Works
@@
Figure 1: SOFAI-v2 architecture with real-time metacognitive governance (supervisory control over fast/automatic and slow/controlled processing)
@@
This section reviews the evolution of solvers for graph coloring, the application of LLMs in sequential decision-making tasks, and the emergence of neurosymbolic approaches for enhancing decision-making processes. We also situate SOFAI-v2 within cognitive control theories that distinguish routine (automatic) processing from controlled, supervised processing (Norman & Shallice, 1986; Cooper & Shallice, 2000).
3.1 Solvers for Graph Coloring
@@
Traditional symbolic solvers have been extensively developed for this problem, leveraging algorithms that systematically explore the search space. Early works like (Brélaz, 1979) introduced efficient heuristics such as the DSATUR algorithm, significantly improving practical performance. (Chaitin, 1982) applied graph coloring to compiler optimization, demonstrating the versatility of symbolic methods. (Golumbic, 2004) provided a comprehensive treatment of graph algorithms, including coloring techniques. Despite their accuracy, symbolic approaches often struggle with scalability as CSP complexity increases, hindered by high computational overhead and slow processing times (Dechter, 2003). Techniques like constraint propagation (Mackworth, 1977) and backtracking algorithms (Bitner & Reingold, 1975) have been employed to enhance efficiency but are limited by the exponential growth of the search space.
DSATUR selects at each step the uncolored vertex with the highest saturation degree (the number of distinct colors among its colored neighbors), breaking ties by degree (Brélaz, 1979). This ordering aggressively focuses search on the most constrained vertices, yielding strong practical performance. With appropriate data structures, per-step selection can be done in near $O(logn)$ update time, but the overall search with backtracking remains exponential in the worst case due to the NP-completeness of graph coloring. In our setting, we couple DSATUR with backtracking under a fixed color budget $k$: the procedure attempts to extend a partial assignment using at most $k$ colors; failure to extend yields a certified UNSAT outcome, while success produces a constructive $k$-coloring witness.
To address these limitations, neural solvers have been explored, utilizing machine learning techniques to approximate solutions for graph coloring. For instance, (Li et al., 2018) proposed using graph neural networks to tackle combinatorial optimization problems. Similarly, (Khalil et al., 2017) introduced a framework that learns heuristics directly from data, improving scalability. However, neural approaches often face challenges in strictly adhering to the constraints inherent in CSPs. In contrast, our SOFAI setup directly targets the decision form of graph coloring with an explicit $k$, integrating DSATUR+backtracking as S2 and using S1/SOFAI components to propose assignments that are subsequently verified. Many end-to-end neural solvers are not readily adaptable to this satisfiability variant, which limits their comparability to our setting; we therefore retain them here primarily for historical context and contrast, while focusing our technical discussion on DSATUR and its role in our pipeline.
3.2 LLMs for Reasoning Tasks
@@
Recent developments in LLMs have opened new avenues for addressing sequential decision-making tasks, including CSPs. Various prompt engineering techniques have been proposed to enhance the reasoning capabilities of LLMs (Brown, 2020; Wei et al., 2022; Kojima et al., 2022; Nye et al., 2021; Wang et al., 2022). (Han et al., 2023) demonstrated that LLMs can arrive at correct answers for graph-based problems through iterative prompting. However, the reasoning capabilities of LLMs have been found to be fundamentally approximate retrieval in nature Kambhampati (2024), limiting their effectiveness in precise reasoning tasks. This limitation is evidenced by studies showing LLMs’ inability to solve planning problems (Valmeekam et al., 2022) and graph coloring problems (Stechly et al., 2024), where strict adherence to constraints is required (Jiang et al., 2023). In cognitive-systems terms, LLM-only routines resemble contention among habitual schemas without sufficient supervisory control (Norman & Shallice, 1986).
3.3 Neurosymbolic Approaches for Reasoning Tasks
@@
Neurosymbolic approaches have gained considerable attention in addressing complex sequential decision-making tasks such as automated planning (Fabiano et al., 2023), constrained grid navigation (Ganapini et al., 2022), and puzzle solving (Lin et al., 2024). These studies demonstrate that integrating neural methods, such as LLMs, with symbolic solvers enhances performance on sequential decision-making tasks. To the best of our knowledge, neurosymbolic techniques have not been explored in the context of graph coloring problems. To our knowledge, prior neurosymbolic studies have not targeted the graph coloring decision problem, nor integrated meta-level feedback with a decision-oriented S2; SOFAI-v2 explicitly addresses this gap. Conceptually, SOFAI-v2 instantiates a neurosymbolic cognitive architecture in which S1 (neural) handles routine proposals while MC+S2 provide supervisory/controlled processing aligned with Norman–Shallice-style accounts (Norman & Shallice, 1986; Cooper & Shallice, 2000).
4 The SOFAI Architecture
@@
The SOFAI architecture, also referred to as SOFAI-v1 here (Booch et al., 2021), leverages a rule-based metacognitive control mechanism that dynamically chooses between S1 and S2 solvers based on a confidence threshold associated with S1’s outputs. S1 solvers are experience-based (and usually data-driven) solvers, while S2 solvers are deliberative (and usually symbolic and rule-based) solvers. SOFAI-v1 has been instantiated to planning (Fabiano et al., 2023) and constrained-grid navigation (Ganapini et al., 2022), showing improvement on the performance compared to either symbolic or data-driven solvers for the same class of problems.
@@
In SOFAI-v1, the metacognitive component only chooses between S1 and S2 solvers, but does not exploit the possible collaboration between these two kinds of solvers, nor does it provide feedback to the solvers if they cannot solve the given problem instance. In this paper we consider a generalized version of this architecture, called SOFAI-v2, shown in Figure 1, where the metacognitive governance (called MC) can provide feedback and examples to a failing S1 solver (that for this paper is an LLM), calling the S1 solver more than once, until either a correct solution is returned or a maximum number of iterations is reached, which triggers the activation of an S2 solver. This implements a neurosymbolic control loop: S1 produces routine proposals (automatic), MC performs verification and trend monitoring (supervisory attention), and S2 executes controlled search when conflicts persist, echoing contention scheduling with a supervisory system (Norman & Shallice, 1986; Cooper & Shallice, 2000).
@@
Formally, let $f_{\text{S1}}(x,M)$ denote the solution produced by S1 for a problem instance $x$ with access to an episodic memory $M$. Here, $M$ consists of previously encountered problem-solution pairs generated by SOFAI-v2, i.e., pairs $(x_{j},f_{\text{SOFAI-v2}}(x_{j}))$, where $f_{\text{SOFAI-v2}}(x_{j})$ represents the final solution that could have been generated by either S1 or S2. In the context of this paper, $x$ corresponds to a graph coloring decision problem instance defined by a graph $G$ and a color bound $k$.
@@
where $x_{j}\in M$ and $\alpha$ is a predefined similarity threshold. This subset, $M_{x}$, provides S1 with additional contextual information from past solutions, enabling it to leverage SOFAI-v2’s previous outputs for experience-guided context.
The correctness of solution for the graph coloring decision problem can be checked and evaluated in polynomial time by MC. The correctness score used by MC for graph coloring solutions is:
$C(f_{S1}(x,M_{x}))=\frac{\sum_{(v,u)\in E}\mathbf{1}[f(v)\neq f(u)]}{|E|}$
where:
  • $f(v)$ denotes the assigned color of vertex $v$,
  • $E$ denotes the set of edges in the graph.
If this correctness meets or exceeds a threshold $\theta_{\text{S1}}$, solution is accepted. Otherwise, MC iteratively provides structured feedback $\mathcal{F}(f_{S1}(x,M_{x}))$ or simplified examples $\mathcal{E}(x)$. We set the correctness threshold to 1, since 1 indicates full constraint adherence for the graph-coloring decision task; this ensures S1 outputs are only accepted when a complete, validator-checked coloring is produced. MC also monitors S1’s solution correctness across iterative feedbacks, identifying improvement trends defined by:
$\forall i\in\{1,\dots,m-1\},\quad C_{i+1}>C_{i}$
Failure to show improvement within $m$ iterations results in MC invoking the fallback S2 solver, ensuring robustness. Thus, the iterative feedback mechanism and episodic memory advances SOFAI-v2 over SOFAI-v1, enhancing solver effectiveness via structured feedback and experience-based learning. Our theoretical commitments are: (i) dual-solver with metacognitive governance, (ii) episodic memory for similarity-based retrieval, and (iii) verify-and-feedback loops grounded in polynomial-time checks. The concrete instantiation (Mistral-7B as S1, DSATUR with backtracking as S2) is an implementation detail; same principles transfer to other domains. Detailed algorithm for the architecture is in Supplementary Material Section 8.
5 Experimental Setup and Results
This section details methodology and experimental results evaluating various solver configurations on graph coloring problems, emphasizing efficiency and success rate under systematic constraints. We also report how iterative feedback and selective escalation can be interpreted as supervisory control improving routine policy performance under increasing constraint conflict (Norman & Shallice, 1986).
5.1 Data Generation and Problem Classification